# Tunable quasi-discrete spectrum of spin waves excited by periodic laser patterns

### v1

@@ Abstract
We present a concept for selective excitation of magnetostatic surface waves with quasi-discrete spectrum using spatially patterned femtosecond laser pulses inducing either ultrafast change of magnetic anisotropy or inverse Faraday effect. We micromagnetically simulate excitation of the waves with periodically patterned uni- or bipolar laser impact. Such excitation yields multiple wavepackets propagating with different group velocities, whose dispersion corresponds to the set of quasi-discrete points. In addition, we show that the frequency of the spectral peaks can be controlled by the polarity of the periodic impact and its spatial period. Presented consideration of multiple spatially periodic SW sources as a whole enables implementation of a comprehensive toolkit of spatio-temporal optical methods for tunable SW excitation and control of its parameters.
@@ II.1 Geometry of MSSW laser excitation
To construct the model of spatially periodic laser excitation of MSSW, we consider bismuth-doped yttrium iron garnet Bi:YIG as a magnetic material, since it supports CMA and IFE for excitation of magnetization dynamics Shelukhin et al. (2018). As we are interested only in the excitation of MSSW, we reduce the model to one-dimensional (1D) geometry (Fig. 1 a), and consider a film, where plane MSSW are excited by a laser spot limited in the $x$ direction and infinitely long in the $y$ direction, as in experiments with elongated pump laser spots Satoh et al. (2012); Kainuma, Matsumoto, and Satoh (2021); Matsumoto et al. (2020); Hioki, Hashimoto, and Saitoh (2020, 2022); Yoshimine et al. (2017). In-plane external magnetic field $\textbf{H}_{\text{ext}}$ is oriented along the $y$ axis, and MSSW propagate perpendicularly to $\textbf{H}_{\text{ext}}$, i.e. along $x$ axis. In our model, the uniaxial magnetic anisotropy axis with the equilibrium parameter $K_{u}$ is oriented along $x$ axis to fulfill the condition of optical excitation of MSSW due to CMA Khokhlov et al. (2019); Filatov et al. (2020). Also, we assume the excitation to be uniform across thickness of the film to exclude the exchange standing spin waves (SSW).
@@ II.2 Physical impacts of the laser pulse
The approximation of the excitation being homogeneous across the thickness is valid for iron garnets with thicknesses ranging from tenths to hundreds of nanometers. In general, however, pump pulse impact in iron garnet films is non-uniform across the thickness in accordance with light penetration depth. This should cause the excitation of SSW or higher-order modes of MSSW. We omit this from the consideration in order to focus on propagation of lowest-order MSSW. We note that high-quality ultrathin garnet films where our assumption is valid are now becoming available Levy et al. (2019); Hartmann et al. (2024).
@@ II.3 Analysis of wavenumber distribution
We calculate the corresponding Fouirer amplitude distribution of the modes as the dependence on the spatial period of the pump pattern $\delta$ (Fig. 1 b and c). In the case of unipolar impact, the distribution has maximal amplitudes around even order modes starting at $k=0$ (Fig. 1 b). In contrast, the bipolar impact leads to emergence of odd order modes (Fig. 1 c). This distribution is multiplied by the overall decrease in amplitude with $k$ governed by the spatial width $\sigma$ of the individual laser spot in the pattern and described by the exponential factor in Eq. (6). The spacing between the spots $\delta$ controls both the wavevnumbers and the amplitudes of the excited modes, as seen in Fig. 1 b and c. For example, increasing $\delta$ leads to excitation of higher-order modes, as their wavenumbers decrease and they fall within a range of $k$ limited by the size of the single spot. Thus, while the total width of the excited wavenumber is determined by $\sigma$, the periodicity establishes the quasi-discrete set of wavenumbers with spectral positions controlled by the spatial period $\delta$. As expected, the quasi-discrete spectrum emerges if $\delta$ exceeds $3\sigma$, and the excitation pattern acquires a contrast higher than $30\%$. We also note that the change of the total number of Gaussians in periodic pattern affects their spectral width (see Appendix B).
@@ III.1 Modeling parameters
Micromagnetic validation of the proposed model is performed using Ubermag Beg, Lang, and Fangohr (2022) with object oriented micromagnetic framework (OOMMF) Donahue and Porter (1999) as a micromagnetic calculator. In simulations, we use the material parameters, typical for Bi:YIG films Zeuschner et al. (2022): magnetization saturation $\mu_{0}M_{s}=175$ mT, exchange energy constant $A_{ex}=3.5\times 10^{-12}$ J/m, Gilbert damping $\alpha=0.01$, and uniaxial anisotropy parameter $K_{u}=5750$ J/m3. The in-plane external magnetic field $\mu_{0}H_{\text{ext}}=100$ mT is applied with a tilt of 1 with respect to $Y$ axis. We set the lateral size of the computational cell as $10\times 10$ nm2 to be smaller than the magnetostatic exchange $\sqrt{2A/(\mu_{0}M_{s}^{2})}=16.9$ nm and the magnetocrystalline exchange lengths $\sqrt{A/K_{u}}=24.7$ nm both Abo et al. (2013). The thickness of the film is 500 nm. The corresponding sagittal cell size is 500 nm to exclude the excitation of SSW and higher-order MSSW. 1D periodic boundary conditions are applied along $Y$ axis. In the simulations, time-domain dynamics is tracked by solving the Landau-Lifshitz-Gilbert equation with the set of energy terms: Zeeman energy, exchange energy, energy of the uniaxial magnetic anisotropy, and magnetostatic energy. CMA is simulated by introducing the spatially non-uniform parameter $\widetilde{K_{u}}(x,t,N,\delta)$ described by (4). IFE is simulated by adding to the system the Zeeman energy term with the transient field in the form (5). The amplitudes of CMA and IFE are close to the experimentally achievable values: $\Delta K_{u}=0.3$ Carpene et al. (2010); Gerevenkov et al. (2021) and $\mu_{0}H_{\text{IFE}}=\pm 1$ T Kimel et al. (2005); Kalashnikova, Kimel, and Pisarev (2015); Kozhaev et al. (2018); Shelukhin et al. (2018). The ground state of the magnetization is calculated as a relaxed state after the initializing of the system with the equilibrium parameters, minimization of its energy, and time domain calculations during 10 ns with an increased Gilbert damping of $\alpha=1$ without excitation.
@@ III.2 Single pulse excitation
First, we consider the excitation of MSSW using a single source, i.e., at $N=0$ in Eq. (2). In the spatio-temporal domain, it is common to monitor the out-of-plane magnetization component $M_{z}$, following the optical pump-probe experiments using Faraday Satoh et al. (2012); Chernov et al. (2017) or polar Kerr Khokhlov et al. (2019); Filatov et al. (2022); Au et al. (2013b) effects. CMA and IFE are used as excitation mechanisms, and the corresponding spatial excitation profiles at $t=t_{0}$ are shown in the left panels of Fig. 2 a and e. The right panels show the corresponding spatio-temporal dependencies of $M_{z}(t,x)$. Within the excited area ($|x|<\sigma$), the laser-induced magnetization precession is observed at $t>t_{0}=25$ ps. The propagation of the MSSW is observed away from the excitation area in the form of a wavepacket. Regardless of the excitation mechanism, the wave phase and group velocities of the same sign are characteristic for MSSW, and waves propagate over a distance of more than 25 $\mu$m within the modeled time window of 4.5 ns (Fig. 2 b and f). In addition, the initial phase of the waves is defined by the circular polarization direction in the case of IFE Kimel et al. (2005); Satoh et al. (2012),
@@
However, upon closer examination, the differences in the dynamics of $M_{z}(t,x)$ are observed between the two excitation mechanisms (right panels of Fig. 2 a and e): a sharper phase front evolution around $x=0$ in the IFE case indicates differences in frequencies between the two mechanisms. To explain it, we consider the FFT of $M_{z}(t)$ traces at various distances from the excitation area (Fig. 2 c and g). At $x=0$ (red lines), an excitation of a wide frequency range is observed with two peacks near 2 and 4 GHz for both IFE and CMA. The low-frequency peak is more pronounced in the IFE case, and the high-frequency peak has a higher frequency in the CMA case (marked with a dashed line). Out of the excitation area (blue and green lines in Fig. 2 c and g), the frequency spectra are similar for the two excitation methods and lie between the frequencies of the initial peaks. The wave dispersions for the two excitation mechanisms obtained by 2D FFT (Fig. 2 d and h) indicate that the two peaks correspond to the minimum and maximum frequencies of MSSW.
@@
The observed differences in the excited spectra are due to the different time dependences for the two excitation methods. In the case of IFE (5) the laser-induced field disappears in a time much shorter than the characteristic period of magnetization dynamics. After pumping, the excited dynamics at $x=0$ has a certain frequency spectrum and then MSSW propagates out of the excitation area with group velocities corresponding to its dispersion. In the case of CMA Eq. (4) the material parameter $K_{u}$ experiences a long-lived partial quenching in the pump area. As condition $\mu_{0}H_{\text{ext}}>2K_{u}/M_{s}$ is satisfied, CMA shifts the dispersion curve to higher frequencies within the excitation area Filatov et al. (2022, 2020); Khokhlov et al. (2019). As a result, the state with minimum frequency inside the excited area corresponds to the states with non-zero group velocity outside this area. This leads to a less pronounced low-frequency peak in the excitation area in the case of CMA (red line in Fig. 2 c). In contrast, the high-frequency part experiences a magnon ”trap” due to CMA, since there are no magnon states at these frequencies outside the pumped area. It leads to an increase in the maximum excited frequencies in the case of CMA and a more pronounced component with the wavenumbers $|k|>5$ rad/$\mu$m observed inside the excitation area (Fig. 2 d). Thus, the differences in the spectra are governed by the parts of the dispersion with near-zero group velocity and are the most pronounced near the excitation area (see Appendix A for details). The propagating MSSW outside the pumped area can be considered identical for CMA and IFE.
@@ III.3 Periodically patterned excitation of MSSW
In all three cases, right after the excitation, the phase front is periodic in space and follows the excitation pattern. In the case of IFE, such a spatially periodic phase is also observed at longer time delays because the material preserves its homogeneous magnetic characteristics. The apparent change of the phase front at longer delay times is a result of constructive and destructive interference of the MSSW. When CMA excitation is considered, there is an evident evolution of the phase front already after 1 temporal period. As seen in the right panel of Fig. 3 a, the phase front within the excitation spots and between them becomes inhomogeneous. Furthermore, the temporal period of the oscillations of $M_{z}$ changes with time (see the temporal evolution at $x=0,\pm 3,\pm 6$$\mu$m). This intricate behavior can be understood as an evolution of the spin-wave distribution inside a set of magnon resonators, which eventually results in a establishing of MSSW dispersion characteristic to a magnonic crystal.
The distribution of $M_{z}(t,x)$ throughout the spatio-temporal area demonstrates the propagation of multiple MSSW wavepackets with different group velocities (Fig. 3 b,d and f). The group velocities of the wavepackets are approximately the same when excited by CMA and unipolar IFE (Fig. 3 b and d). The change in IFE polarity results in the excitation of MSSW wavepackets with decreased group velocities (Fig. 3 d and f). Although the temporal period of magnetization dynamics appeared to be the same within the excitation area for the uni- and bipolar IFE, the temporal period of the outgoing wavepackets decreases when the bipolar IFE is simulated.
@@ IV Conclusion
The developed model proves and extends the flexible reconfigurability offered by optical methods of SW excitation. Indeed, tunable periodical patterning of the spatial profile of the laser pulse can be achieved, for example, in the double pump geometry used for the creation of transient gratings Janušonis et al. (2016); Carrara et al. (2022) or by a spatial light modulator Kolosvetov et al. (2022); Kolosvetov and Chernov (2023). The model may find practical implementation in the development and modeling of complicated magnonic systems with multiple SW sources, which is one of the required conditions for the design of magnon logic devices Talmelli et al. (2020). Also, the presented approach may be useful to describe magnonic neuromorphic computing devices, as it offers a consideration of the system with a number of SW sources. We note that our model can be readily improved to include in the consideration 2- and 3-dimensional spatially patterned SW sources and to excite and control other types of SW. As for the next steps, we suggest the temporal modulation of the laser impact in magnonic systems, in addition to the spatial patterning considered here, as it makes possible to implement a comprehensive toolkit of spatio-temporal methods for tunable SW excitation and control of its parameters Jäckl et al. (2017); Muralidhar et al. (2020); Hula et al. (2022).
@@ ACKNOWLEDGMENTS
Appendix A Spectral difference between CMA and IFE
Appendix B Wavenumber distribution dependence on number of Gaussians $N$

### v2

@@ Abstract
We present a concept for selective excitation of magnetostatic surface waves with quasi-discrete spectrum using spatially patterned femtosecond laser pulses inducing either ultrafast change of magnetic anisotropy or inverse Faraday effect. We micromagnetically simulate excitation of the waves with periodically patterned uni- or bipolar laser impact. Such excitation yields multiple wavepackets propagating with different group velocities, whose dispersion corresponds to the set of quasi-discrete points. In addition, we show that the frequency of the spectral peaks can be controlled by the polarity of the periodic impact and its spatial period. Presented consideration of multiple spatially periodic magnetostatic surface wave sources as a whole enables implementation of a comprehensive toolkit of spatio-temporal optical methods for tunable excitation and control of spin wave parameters.
@@ II.1 Geometry of MSSW laser excitation
To construct the model of spatially periodic laser excitation of MSSW, we consider bismuth-doped yttrium iron garnet Bi:YIG as a magnetic material, since it supports CMA and IFE for excitation of magnetization dynamics Shelukhin et al. (2018). As we are interested only in the excitation of MSSW, we reduce the model to one-dimensional (1D) geometry (Fig. 1 a), and consider a film, where plane MSSW are excited by a laser spot limited in the $x$ direction and infinitely long in the $y$ direction, as in experiments with elongated pump laser spots Satoh et al. (2012); Kainuma, Matsumoto, and Satoh (2021); Matsumoto et al. (2020); Hioki, Hashimoto, and Saitoh (2020, 2022); Yoshimine et al. (2017). This also excludes an excitation of backward volume magnetostatic spin waves with nonzero wavenumber $k$. In-plane external magnetic field $\textbf{H}_{\text{ext}}$ is oriented along the $y$ axis, and MSSW propagate perpendicularly to $\textbf{H}_{\text{ext}}$, i.e. along $x$ axis. In our model, the uniaxial magnetic anisotropy axis with the equilibrium parameter $K_{u}$ is oriented along $x$ axis to fulfill the condition of optical excitation of MSSW due to CMA Khokhlov et al. (2019); Filatov et al. (2020). Also, we assume the excitation to be uniform across the thickness of the film to exclude the formation of perpendicular standing spin waves (PSSW).
@@ II.2 Physical impacts of the laser pulse
The approximation of the excitation being homogeneous across the thickness is valid for iron garnets with thicknesses ranging from tenths to hundreds of nanometers. In general, however, pump pulse impact in iron garnet films is non-uniform across the thickness in accordance with light penetration depth, which has typical values of about 100 nm for photons with energy above absorption edge Wemple et al. (1974) used for CMA and two-three orders of magnitude larger for photons below absorption edge used for IFE. This should cause the excitation of PSSW. We omit this from the consideration in order to focus on propagation of MSSW. We note that high-quality ultrathin garnet films where our assumption is valid are now becoming available Levy et al. (2019); Hartmann et al. (2024).
@@ II.3 Analysis of wavenumber distribution
We calculate the corresponding Fouirer amplitude distribution of the modes as the dependence on the spatial period of the pump pattern $\delta$ (Fig. 1 b and c). In the case of unipolar impact, the distribution has maximal amplitudes around even order modes starting at $k=0$ (Fig. 1 b). In contrast, the bipolar impact leads to emergence of odd order modes (Fig. 1 c). This distribution is multiplied by the overall decrease in amplitude with $k$ governed by the spatial width $\sigma$ of the individual laser spot in the pattern and described by the exponential factor in Eq. (6). The spacing between the spots $\delta$ controls both the wavevnumbers and the amplitudes of the excited modes, as seen in Fig. 1 b and c. For example, increasing $\delta$ leads to excitation of higher-order modes, as their wavenumbers decrease and they fall within a range of $k$ limited by the size of the single spot. Thus, while the total width of the excited wavenumber is determined by $\sigma$ Satoh et al. (2012); Jäckl et al. (2017), the periodicity establishes the quasi-discrete set of wavenumbers with spectral positions controlled by the spatial period $\delta$. As expected, the quasi-discrete spectrum emerges if $\delta$ exceeds $3\sigma$, and the excitation pattern acquires a contrast higher than $30\%$. We also note that the change of the total number of Gaussians in periodic pattern affects their spectral width (see Appendix A).
@@ III.1 Modeling parameters
Micromagnetic validation of the proposed model is performed using Ubermag Beg, Lang, and Fangohr (2022) with object oriented micromagnetic framework (OOMMF) Donahue and Porter (1999) as a micromagnetic calculator. In simulations, we use the material parameters, typical for Bi:YIG films Zeuschner et al. (2022): magnetization saturation $\mu_{0}M_{s}=175$ mT, exchange energy constant $A_{ex}=3.5\times 10^{-12}$ J/m, Gilbert damping $\alpha=0.01$, and uniaxial anisotropy parameter $K_{u}=5750$ J/m3. The in-plane external magnetic field $\mu_{0}H_{\text{ext}}=100$ mT is applied with a tilt of 1 with respect to $y$ axis. We set the lateral size of the computational cell as $10\times 10$ nm2 to be smaller than the magnetostatic exchange length $\sqrt{2A/(\mu_{0}M_{s}^{2})}=16.9$ nm and the magnetocrystalline exchange length $\sqrt{A/K_{u}}=24.7$ nm Abo et al. (2013). The thickness of the film is 500 nm. The corresponding sagittal cell size 500 nm allows excluding the excitation of PSSW. 1D periodic boundary conditions are applied along $y$ axis. In the simulations, time-domain dynamics is tracked by solving the Landau-Lifshitz-Gilbert equation with the set of energy terms: Zeeman energy, exchange energy, energy of the uniaxial magnetic anisotropy, and magnetostatic energy. CMA is simulated by introducing the spatially non-uniform parameter $\widetilde{K_{u}}(x,t,N,\delta)$ described by (4). IFE is simulated by adding to the system the Zeeman energy term with the transient field in the form (5). The magnitudes of CMA and IFE are close to the experimentally achievable values: $\Delta K_{u}=0.3$ Carpene et al. (2010); Gerevenkov et al. (2021) and $\mu_{0}H_{\text{IFE}}=\pm 1$ T Kimel et al. (2005); Kalashnikova, Kimel, and Pisarev (2015); Kozhaev et al. (2018); Shelukhin et al. (2018). The ground state of the magnetization is calculated as a relaxed state after the initializing of the system with the equilibrium parameters, minimization of its energy, and time domain calculations during 10 ns with an increased Gilbert damping of $\alpha=1$ without excitation. One total individual simulation including numerical Fourier analysis to get the dispersion relations took approximately 1.5 hours using 16-thread CPU at 5 GHz operating frequency.
@@ III.2 Single pulse excitation
[…]
@@
However, upon closer examination, the differences in the dynamics of $M_{z}(t,x)$ are observed between the two excitation mechanisms (right panels of Fig. 2 a and e): a sharper phase front evolution around $x=0$ in the IFE case indicates differences in frequencies between the two mechanisms. To explain it, we consider the FFT of $M_{z}(t)$ traces at various distances from the excitation area (Fig. 2 c and g). At $x=0$ (red lines), an excitation of a wide frequency range is observed with two peacks near 2 and 4 GHz for both IFE and CMA. The low-frequency peak is more pronounced in the IFE case, and the high-frequency peak has a higher frequency in the CMA case (marked with a dashed line). The wave dispersions for the two excitation mechanisms obtained by 2D FFT (Fig. 2 d and h) indicate that the two peaks correspond to the minimum and maximum frequencies of MSSW. This is consistent with the experimental data obtained with a round laser spot in Ref. Chernov et al. (2017) except for the absence of backward volume waves suppressed in our case due to the elongated excitation area shape. Out of the excitation area (blue and green lines in Fig. 2 c and g), the frequency spectra are similar for the two excitation methods and lie between the frequencies of the initial peaks.
@@
Let us discuss the differences between IFE and CMA in the excited spectra at $x=0$$\mu$m. We attribute them to the different time dependencies for the two excitation methods. In the case of IFE (5) the laser-induced field disappears in a time much shorter than the characteristic period of magnetization dynamics. After pumping, the excited dynamics at $x=0$ has a certain frequency spectrum and then MSSW propagates out of the excitation area with group velocities corresponding to its dispersion. Slow waves with frequency around 4 GHz do not leave the excitation area and manifest themselves as a pronounced peak in FFT spectrum at $x=0$. In the case of CMA Eq. (4) the material parameter $K_{u}$ experiences a long-lived partial quenching in the pump area. At magnetization along the hard anisotropy axis, the effective field is $|\mu_{0}H_{\text{ext}}-2K_{u}/M_{s}|$. Decrease of $K_{u}$ leads to increase in the total effective field and hence to increase of the magnetization dynamics frequencies within the excitation area Filatov et al. (2022, 2020); Khokhlov et al. (2019). As a result, the state with minimum frequency (ferromagnetic resonance) inside the excited area corresponds to the states with non-zero group velocity outside this area. Thus, a low-frequency peak in the case of CMA is less pronounced (red line in Fig. 2 c). In contrast, the high-frequency part experiences a magnon ”trap” due to CMA, since there are no magnon states at these frequencies outside the pumped area. It leads to an increase in the maximum excited frequencies in the case of CMA and a more pronounced component with the wavenumbers $|k|>5$ rad/$\mu$m observed inside the excitation area (Fig. 2 d). Thus, the differences in the spectra are governed by the parts of the dispersion with near-zero group velocity and are most pronounced near the excitation area (see Appendix C for details). The propagating MSSW outside the pumped area can be considered identical for CMA and IFE.
@@ III.3 Periodically patterned excitation of MSSW
As in the case of single spot excitation discussed above, magnetization dynamics possesses distinct features within the area subjected to the laser pattern and outside this area depending on the excitation mechanism. First, comparing $M_{z}(t,x)$ within the excitation area, we observe a pronounced difference of phase fronts depending both on the excitation mechanism, CMA or unipolar IFE, and on the polarity of IFE (right panels in Fig. 3 a,c and e). In all three cases, right after the excitation, the phase front is periodic in space and follows the excitation pattern. In the case of IFE, such a spatially periodic phase is also observed at longer time delays because the material preserves its homogeneous magnetic characteristics. The apparent change of the phase front at longer delay times is a result of constructive and destructive interference of the MSSW. When CMA excitation is considered, there is an evident evolution of the phase front already after 1 temporal period. As seen in the right panel of Fig. 3 a, the phase front within the excitation spots and between them becomes inhomogeneous. Furthermore, the temporal period of the oscillations of $M_{z}$ changes with time (see the temporal evolution at $x=0,\pm 3,\pm 6$$\mu$m). This intricate behavior can be understood as an evolution of the spin-wave distribution inside a set of magnon resonators, which eventually results in a establishing of MSSW dispersion characteristic to a magnonic crystal.
The distribution of $M_{z}(t,x)$ throughout the spatio-temporal area demonstrates the propagation of multiple MSSW wavepackets with different group velocities (Fig. 3 b,d and f). The group velocities of the wavepackets are approximately the same when excited by CMA and unipolar IFE (Fig. 3 b and d). The change in IFE polarity results in the excitation of MSSW wavepackets with decreased group velocities (Fig. 3 d and f). Although the temporal period of magnetization dynamics appeared to be the same within the excitation area for the uni- and bipolar IFE, the temporal period of the outgoing wavepackets decreases when the bipolar IFE is simulated. Individual $M_{z}$ time traces at different $x$ positions for all simulated cases are shown in Appendix D.
@@ IV Conclusion
The developed model proves and extends the flexible reconfigurability offered by optical methods of SW excitation. Indeed, tunable periodical patterning of the spatial profile of the laser pulse can be achieved, for example, in the double pump geometry used for the creation of transient gratings Janušonis et al. (2016); Chang et al. (2018); Carrara et al. (2022) or by a spatial light modulator Kolosvetov et al. (2022); Kolosvetov and Chernov (2023). The model may find practical implementation in the development and modeling of complicated magnonic systems with multiple SW sources, which is one of the required conditions for the design of magnon logic devices Talmelli et al. (2020). Also, the presented approach may be useful to describe magnonic neuromorphic computing devices, as it offers a consideration of the system with a number of SW sources. We note that our model can be readily improved to include in the consideration 2- and 3-dimensional spatially patterned SW sources and to excite and control other types of SW. As for the next steps, we suggest the temporal modulation of the laser impact in magnonic systems, in addition to the spatial patterning considered here, as it makes possible to implement a comprehensive toolkit of spatio-temporal methods for tunable SW excitation and control of its parameters Jäckl et al. (2017); Muralidhar et al. (2020); Hula et al. (2022). Furthermore, including elastodynamics and spin-phonon interaction in the model Deb et al. (2021); Azovtsev and Pertsev (2023) would allow accounting for coherent acoustic waves driven by laser-induced thermal transient gratings that offer a promising way to control spin wave parameters due to magnetoelastic coupling Janušonis et al. (2016); Chang et al. (2018); Yaremkevich et al. (2023).
@@ ACKNOWLEDGMENTS
Appendix A Wavenumber distribution dependence on number of Gaussians $N$